# VLM Can Be a Good Assistant: Enhancing Embodied Visual Tracking with Self-Improving Visual-Language Models

### v1

@@ Abstract
We introduce a novel self-improving framework that enhances Embodied Visual Tracking (EVT) with Visual-Language Models (VLMs) to address the limitations of current active visual tracking systems in recovering from tracking failure. Our approach combines the off-the-shelf active tracking methods with VLMs’ reasoning capabilities, deploying a fast visual policy for normal tracking and activating VLM reasoning only upon failure detection. The framework features a memory-augmented self-reflection mechanism that enables the VLM to progressively improve by learning from past experiences, effectively addressing VLMs’ limitations in 3D spatial reasoning. Experimental results demonstrate significant performance improvements, with our framework boosting success rates by $72\%$ with state-of-the-art RL-based approaches and $220\%$ with PID-based methods in challenging environments. This work represents the first integration of VLM-based reasoning to assist EVT agents in proactive failure recovery, offering substantial advances for real-world robotic applications that require continuous target monitoring in dynamic, unstructured environments. Project website: https://sites.google.com/view/evt-recovery-assistant.
@@ I INTRODUCTION
Concurrently, the field has witnessed remarkable advances in Visual-Language Models (VLMs) [25], which have demonstrated exceptional cross-domain generalization capabilities and powerful scene-reasoning abilities. However, these models have notable limitations in robotics applications. Most VLMs are primarily trained on static images or passive video understanding without action capabilities, resulting in limited spatial reasoning [2] and navigation planning [24] in complex 3D environments. Despite these constraints, recent VLMs like GPT-4V have shown impressive zero-shot performance and in-context learning ability in understanding complex visual scenes, which presents a promising direction for addressing the limitations of current EVT approaches.
@@
To bridge this gap, we propose a novel self-improving framework that integrates a Visual-Language Model (VLM) to assist EVT agents in recovering from tracking failures. Our approach combines the off-the-shelf traditional tracking algorithms with the powerful reasoning capabilities of VLMs, creating a complementary system that leverages each component’s strengths while mitigating their individual weaknesses. Specifically, we employ an RL-based visual tracking policy [29] to follow the target in nominal conditions and only activate the more computationally intensive VLM reasoning when failures are detected. The system identifies target loss through a reliable segmentation-based detection mechanism that continuously monitors the target’s presence. Upon failure detection, the VLM analyzes the historical observation sequence leading to the failure, interprets the environmental context, and generates a structured sequence of recovery actions as movement suggestions. Most importantly, to overcome the inherent limitations of VLMs in 3D spatial reasoning and to optimize recovery strategies for unknown failure cases, we introduce a memory-augmented self-reflection mechanism that enables the VLM to progressively self-improve by analyzing, storing, and learning from its historical experiences. This approach creates a continuously evolving recovery capability that becomes increasingly effective at handling complex tracking failures over time.
@@ II-A Embodied Visual Tracking in Complex Environments
Visual tracking is a fundamental capability for mobile robots operating in dynamic environments, requiring systems to actively control camera views and maintain awareness of surroundings while following targets [6, 7]. Previous research in this domain has approached the challenge from several perspectives. Traditional approaches relied primarily on geometric methods [4, 18] and filter-based tracking techniques [21] to maintain target awareness. While effective in controlled environments, these methods often struggled with occlusions and dynamic scenes. Building on these foundations, specialized person-following robots [14, 19] emerged, predictive modeling to maintain line-of-sight tracking. More recently, reinforcement learning (RL) based approaches have made substantial advances in tracking robustness [17, 31]. Researchers further employ spatial-temporal neural networks and adversarial mechanisms to handle short-term occlusions and distractions  [28, 27]. Advanced state representation and RL strategies  [29] have further improved training efficiency and enabled short-term trajectory prediction for brief occlusion handling. Despite these improvements, existing robotic systems still struggle to proactively recover from significant tracking failures, particularly in scenarios involving complex environmental structures. In this work, we propose a self-improving framework that integrates a Visual-Language Model (VLM) to assist visual tracking agents in recovering from such failures.
@@ II-B Vision-Language Model for Embodied Vision
Figure 2: The framework for integrating Visual-Language Models (VLMs) with active tracking policies. The framework follows a structured recovery approach when target tracking fails, consisting of five main steps: (1) Failure Detection that uses segmentation-based target identification; (2) Failure Case Analysis through chain-of-thought reasoning to understand the environmental context; (3) Movement Suggestions that provide structured action plans with directional, environmental, and conditional triggers; and (4) Memory-Augmented Self-Reflection that plans recovery sequences and optimizes actions based on stored experiences. (5)Reflection Insights Generation that summarizes the entire recovery plan and gives an adjustment suggestion when the plan fails. The framework enables agents to recover from occlusions and extended target loss by leveraging the reasoning capabilities of VLMs, e.g., GPT-4o, memory management for historical context, self-reflection for continuous improvement, and a robust visual tracking policy for sustained tracking once recovery is achieved.
@@ IV-A Visual Tracking Policy
IV-B Visual-Language Reasoning for Failure Recovery
@@ V-A Experiment Setup
Baseline. We compare our approach against four relevant baselines. 1) PID: A classical control approach that calculates the IoU between the detected target bounding box and a desired bounding box position, using a PID controller to adjust agent actions to maximize this IoU. 2) OpenVLA [12]: While primarily designed for robotic manipulation, we fine-tuned this popular model on our collected dataset to adapt it for tracking tasks, providing a comparison with a recent domain-adjacent approach. 3) GPT-4o: A direct application of the state-of-the-art visual-language model using the same action space as our framework’s reasoning module. This baseline demonstrates the raw capabilities of advanced VLMs without our specialized recovery framework. 4) SOTA RL [29]: The state-of-the-art RL-based method that integrates the visual foundation model and offline reinforcement learning to learn a generalizable tracking policy. We follow the official repository’s data collection and training procedures, using 50k-step trajectories collected in a room with domain randomization for training.

### v2

# VLM Can Be a Good Assistant: Enhancing Embodied Visual Tracking with Self-Improving Vision-Language Models

@@ Abstract
We introduce a novel self-improving framework that enhances Embodied Visual Tracking (EVT) with Vision-Language Models (VLMs) to address the limitations of current active visual tracking systems in recovering from tracking failure. Our approach combines the off-the-shelf active tracking methods with VLMs’ reasoning capabilities, deploying a fast visual policy for normal tracking and activating VLM reasoning only upon failure detection. The framework features a memory-augmented self-reflection mechanism that enables the VLM to progressively improve by learning from past experiences, effectively addressing VLMs’ limitations in 3D spatial reasoning. Experimental results demonstrate significant performance improvements, with our framework boosting success rates by $72\%$ with state-of-the-art RL-based approaches and $220\%$ with PID-based methods in challenging environments. This work represents the first integration of VLM-based reasoning to assist EVT agents in proactive failure recovery, offering substantial advances for real-world robotic applications that require continuous target monitoring in dynamic, unstructured environments. Project website: https://sites.google.com/view/evt-recovery-assistant.
@@ I INTRODUCTION
Concurrently, the field has witnessed remarkable advances in Vision-Language Models (VLMs) [25], which have demonstrated exceptional cross-domain generalization capabilities and powerful scene-reasoning abilities. However, these models have notable limitations in robotics applications. Most VLMs are primarily trained on static images or passive video understanding without action capabilities, resulting in limited spatial reasoning [2] and navigation planning [24] in complex 3D environments. Despite these constraints, recent VLMs like GPT-4V have shown impressive zero-shot performance and in-context learning ability in understanding complex visual scenes, which presents a promising direction for addressing the limitations of current EVT approaches.
@@
To bridge this gap, we propose a novel self-improving framework that integrates a Vision-Language Model (VLM) to assist EVT agents in recovering from tracking failures. Our approach combines the off-the-shelf traditional tracking algorithms with the powerful reasoning capabilities of VLMs, creating a complementary system that leverages each component’s strengths while mitigating their individual weaknesses. Specifically, we employ an RL-based visual tracking policy [29] to follow the target in nominal conditions and only activate the more computationally intensive VLM reasoning when failures are detected. The system identifies target loss through a reliable segmentation-based detection mechanism that continuously monitors the target’s presence. Upon failure detection, the VLM analyzes the historical observation sequence leading to the failure, interprets the environmental context, and generates a structured sequence of recovery actions as movement suggestions. Most importantly, to overcome the inherent limitations of VLMs in 3D spatial reasoning and to optimize recovery strategies for unknown failure cases, we introduce a memory-augmented self-reflection mechanism that enables the VLM to progressively self-improve by analyzing, storing, and learning from its historical experiences. This approach creates a continuously evolving recovery capability that becomes increasingly effective at handling complex tracking failures over time.
@@ II-A Embodied Visual Tracking in Complex Environments
Visual tracking is a fundamental capability for mobile robots operating in dynamic environments, requiring systems to actively control camera views and maintain awareness of surroundings while following targets [6, 7]. Previous research in this domain has approached the challenge from several perspectives. Traditional approaches relied primarily on geometric methods [4, 18] and filter-based tracking techniques [21] to maintain target awareness. While effective in controlled environments, these methods often struggled with occlusions and dynamic scenes. Building on these foundations, specialized person-following robots [14, 19] emerged, predictive modeling to maintain line-of-sight tracking. More recently, reinforcement learning (RL) based approaches have made substantial advances in tracking robustness [17, 31]. Researchers further employ spatial-temporal neural networks and adversarial mechanisms to handle short-term occlusions and distractions  [28, 27]. Advanced state representation and RL strategies  [29] have further improved training efficiency and enabled short-term trajectory prediction for brief occlusion handling. Despite these improvements, existing robotic systems still struggle to proactively recover from significant tracking failures, particularly in scenarios involving complex environmental structures. In this work, we propose a self-improving framework that integrates a Vision-Language Model (VLM) to assist visual tracking agents in recovering from such failures.
@@ II-B Vision-Language Model for Embodied Vision
Figure 2: The framework for integrating Vision-Language Models (VLMs) with active tracking policies. The framework follows a structured recovery approach when target tracking fails, consisting of five main steps: (1) Failure Detection that uses segmentation-based target identification; (2) Failure Case Analysis through chain-of-thought reasoning to understand the environmental context; (3) Movement Suggestions that provide structured action plans with directional, environmental, and conditional triggers; and (4) Memory-Augmented Self-Reflection that plans recovery sequences and optimizes actions based on stored experiences. (5)Reflection Insights Generation that summarizes the entire recovery plan and gives an adjustment suggestion when the plan fails. The framework enables agents to recover from occlusions and extended target loss by leveraging the reasoning capabilities of VLMs, e.g., GPT-4o, memory management for historical context, self-reflection for continuous improvement, and a robust visual tracking policy for sustained tracking once recovery is achieved.
@@ IV-A Visual Tracking Policy
IV-B Vision-Language Reasoning for Failure Recovery
@@ V-A Experiment Setup
Baseline. We compare our approach against four relevant baselines. 1) PID: A classical control approach that calculates the IoU between the detected target bounding box and a desired bounding box position, using a PID controller to adjust agent actions to maximize this IoU. 2) OpenVLA [12]: While primarily designed for robotic manipulation, we fine-tuned this popular model on our collected dataset to adapt it for tracking tasks, providing a comparison with a recent domain-adjacent approach. 3) GPT-4o: A direct application of the state-of-the-art vision-language model using the same action space as our framework’s reasoning module. This baseline demonstrates the raw capabilities of advanced VLMs without our specialized recovery framework. 4) SOTA RL [29]: The state-of-the-art RL-based method that integrates the visual foundation model and offline reinforcement learning to learn a generalizable tracking policy. We follow the official repository’s data collection and training procedures, using 50k-step trajectories collected in a room with domain randomization for training.